# Multimodal Crash Likelihood Prediction: A Complexity-Infused Approach Integrating Semantic, Contextual, and Driving Features

### v1

@@ Abstract
Predicting crash likelihood in complex driving environments is essential for improving traffic safety and advancing autonomous driving. Previous studies have used statistical models and deep learning to predict crashes based on semantic, contextual, or driving features, but none have examined the combined influence of these factors, termed “roadway complexity” in this study. This paper introduces a two-stage framework that integrates roadway complexity features for crash prediction. In the first stage, an encoder extracts hidden contextual information from these features, generating complexity-infused features. The second stage uses both original and complexity-infused features to predict crash likelihood, achieving an accuracy of 87.98% with original features alone and 90.15% with the added complexity-infused features. Ablation studies confirm that a combination of semantic, driving, and contextual features yields the best results, which emphasize their role in capturing roadway complexity. Additionally, complexity index annotations generated by Large Language Models outperform those by Amazon Mechanical Turk, highlighting the potential of automated tools for accurate, scalable crash prediction systems.
@@ 1 Introduction
Figure 1: The model structure. The model takes raw images as input and generates semantic, driving, and contextual features from various sources to predict the crash likelihood as the output. It consists of an encoder that learns hidden features from the semantic, driving, and contextual data, which are infused with the complexity index. The prediction model then utilizes all the available features, including the complexity-infused features, to predict the crash likelihood. Example data is shown above each feature source.
Understanding factors contributing to crashes is essential for improving traffic safety and advancing autonomous vehicle design. A 2015 NHTSA report [20] found that 94% of the 5,470 crashes from 2005 to 2007 were due to human errors like inattention, distractions, excessive speed, and misjudgment.
Research has sought to predict crash frequency and rates based on environmental factors and driver behavior. Statistical models have linked crash occurrences to variables such as speed, traffic volume, weather, and road design [22, 19, 24, 6, 23], while recent studies have used deep learning to analyze real-time images and sensor data from vehicles and infrastructure to detect hazards [11, 8]. Though promising, these models rely on either imagery or behavioral data, without integrating both to predict crashes.
@@
Driving behavior, such as speed adjustments in response to poor visibility or narrow lanes, is also influenced by scene complexity. Speed and acceleration patterns adjust based on obstacles and conditions [12]. Integrating behavior data with scene information deepens our understanding of driver interactions with their environment, improving crash risk modeling. Thus, we define roadway complexity as the combination of scene complexity and driving behavior.
Extracting hidden context from this combined data is essential. Previous studies have shown that fusing situational and memory-based features [27], as well as road graph and motion history data [16], enhances situation awareness and motion prediction, respectively. Building on this, we incorporate feature fusion to capture both explicit and implicit features of roadway scenes.
To address the challenges of (1) studying roadway complexity more holistically by incorporating both imagery and behavioral data, and (2) investigating the direct and indirect relationships between roadway complexity and crashes, we propose a two-stage fusion prediction model framework. This framework learns hidden features from the fusion of scene-level semantic, driving-level, and contextual features and uses the combination of these features to estimate crash likelihood, as illustrated in Fig. 1. Our model, which is trained on a naturalistic dataset and historical crash data, consists of an encoder that captures the hidden context of the roadway complexity and a prediction module that uses these features to estimate crash likelihood. The contributions of this paper are:
@@
    We take a holistic approach to studying roadway complexity by considering both imagery information, i.e., specifically, scene-level semantic and contextual features, and driving behavioral data, and we empirically demonstrate the necessity of incorporating all these elements.
@@
    We argue that crashes can be directly and/or indirectly associated with roadway complexity. We introduce a two-stage framework that extracts hidden contextual features of roadway scene complexity by fusing scene-level, driving-level, and contextual data, and we investigate the predictive power of both direct and hidden features on crash likelihood.
@@
    We compare the complexity index annotations from Amazon Mechanical Turk and Large Language Models (LLMs) in terms of their capability to predict crash likelihood and find that LLMs-generated annotations consistently exhibited better predictive performance. This can enhance the development of real-time crash prediction systems and inspire the integration of automated annotation tools for improved accuracy and scalability.
@@ 2 Related Work
There has been extensive research on understanding roadway scenes using both Convolutional Neural Network (CNN)-based and transformer-based methods. Panoptic segmentation, which unifies semantic and instance segmentation, offers a holistic understanding by labeling all pixels and differentiating object instances, making it especially useful in structured environments like roadways [13]. The task has been successfully applied to driving scene datasets, such as CityScapes [2], which provides high-resolution images for semantic urban scene understanding. Recent advancements in transformer-based models have further improved performance on scene understanding tasks. Among these, OneFormer [9] has achieved state-of-the-art panoptic quality scores [13] while maintaining computational efficiency, making it suitable for real-time scene understanding in autonomous driving. Given its strong performance, the OneFormer model was utilized in this study to extract scene-level visual features from the driving environments.
Understanding the driving context is essential for both human drivers and autonomous systems, as it involves not only the perception of roadway scene elements but also can help interpret the interactions between these elements. Recent advancements in LLMs have demonstrated their potential to go beyond pixel-level interpretation, incorporating higher-level reasoning into driving tasks. LLMs have been applied to extract contextual information from driving environments by mapping raw sensory data (e.g., images, LiDAR) to higher-level contextual descriptions [21]. This is particularly useful for driving research, where understanding the “intent” of other road users, predicting possible hazards, or inferring rules of the road from visual cues can significantly enhance safety and performance [3, 26]. The application of LLMs indicates a leap forward as they can go beyond object detection to understand the context of the driving scenarios.
@@ 3 Methods
In this section, we outline the dataset that was used in the study and the methodology for generating the features used in the crash likelihood prediction model. We detail the architecture of the complexity-infused encoder and describe the structure of the final prediction model.
@@ 3.1 Dataset Overview
This study utilized the MIT-AVT dataset as input [5]. The dataset includes a variety of multimodal data sources, such as 20-second raw video clips from the forward-facing camera, 30 Hz CAN bus data, 30 Hz GPS data, and contextual metadata features, as described by Ding et al. [4].
For the purposes of this research, 500 video clips were selected from the original dataset, categorized as follows: 100 highway scenarios, 100 rural scenarios, 100 urban scenarios, 75 bridge scenarios, 75 overpass scenarios, and 75 crash hotspot scenarios. The methodology for generating crash hotspot data is detailed in Section 3.3.
From each selected video clip, frames were extracted based on a fixed-distance sampling method, with one frame captured for every 20 meters traveled. The extraction began with the first frame of each clip and continued until the distance to the subsequent frame was less than 20 meters. Overall, there are around 10,300 frames extracted.
@@ 3.2 Feature Generation
For each image frame, three sets of features were generated: semantic features derived from a computer vision algorithm, driving features extracted from the CAN bus data, and contextual features generated from the LLMs.
@@ 3.2.1 Semantic Features
The semantic features were generated using the OneFormer algorithm via Hugging Face’s API [^1] https://huggingface.co/docs/transformers/en/model_doc/oneformer. The model outputs pixel-level semantic classifications for various objects such as cars, pedestrians, bicycles, roads, traffic signs, sidewalks, buildings, vegetation, and sky. To better understand the effect of complexity on driver behavior, a lead-car region was defined [25]. An example of the OneFormer algorithm’s output is provided in Fig. 2.
@@
For each image, the percentage of pixels corresponding to each class relative to the total frame size was calculated. Similarly, the percentage of pixels for each class within the lead-car region was computed. Additionally, the number of cars, pedestrians, buses, bicycles, and motorcycles was counted in both the full frame and the lead-car regions. This process resulted in 50 initial semantic features. After closer examination, features with minimal variability (more than 90% being 0) were removed, reducing the final set to 17 semantic features. The final set of features is coded as:
@@
3.2.2 Driving Features
The driving features were extracted from the CAN bus data. For each frame and the corresponding 20-meter segment, the following 9 features were computed: current speed, mean speed, standard deviation of speed, mean longitudinal acceleration, standard deviation of longitudinal acceleration, minimum longitudinal acceleration, maximum longitudinal acceleration, raw deviation from the speed limit, and normalized deviation from the speed limit. The raw and normalized deviation from the speed limit are defined as follows:
@@ 3.2.3 Contextual Features
The contextual features include road characteristics, which were generated using the GPT-4o-2024-08-06 model  [^3] https://platform.openai.com/docs/models/gpt-4o. The prompt used in this step is illustrated in Fig. 3. In the prompt, several questions were asked to gather contextual road characteristics, including information on weather conditions, road conditions, traffic conditions, visibility levels, time of day, road layout, road type, and lane width. Initially, each question was presented in an open-ended format to generate a pool of possible answers. The prompt was then refined based on these responses, providing predefined options to ensure consistency across the answers. The results were output in JSON format. The prompt was run three times on each image to ensure data quality. The answer agreed by the majority will be chosen as the final answer to the questions.
@@ 3.3 Ground Truth Generation
There are two ground truths in this study. As illustrated in Figure 1, for the encoder, the complexity index is collected to quantify the overall complexity level perceived by the driver in a given roadway scene. For the prediction, crash likelihood is computed and serves as the ground truth for estimating crash probability.
@@ 3.3.1 Complexity Index
The complexity index was generated from two sources: AI and humans. For AI, the GPT-4o-2024-08-06 model was used along with the contextual feature generation process, as shown in Fig. 3. In this approach, the model generated a complexity score on a scale from 0 to 10 to describe the complexity and demanding level of the roadway scenes.
The human-generated complexity indices relied on Amazon Mechanical Turk (MTurk) for annotations. The task was designed to assess the complexity level of roadway scenes. Workers were shown image frames and asked to rate the complexity of each scene on a scale from 1 to 10. Only workers with a high approval rating, at least 500 completed tasks, and residing in the US were selected. A pilot study was conducted with 500 images, where 10 workers annotated the same image. The results showed a relatively high level of agreement among workers. Based on this, in the official round, each scene was annotated by 3 workers, and the final complexity score was determined by averaging their responses.
3.3.2 Crash Likelihood
To generate the crash likelihood, historical crash data was obtained from the Massachusetts Department of Transportation’s IMPACT app  [^4] https://apps.impact.dot.state.ma.us/cdp/home. GPS data from crashes over a 5-year period (2018 to 2022) was aggregated. A Kernel Density Estimation (KDE) method was then applied to this GPS data to create a continuous scale representing crash likelihood. The Kernel Density Estimation is formulated as follows:
@@ 3.3.2 Crash Likelihood
After applying the KDE method, each location was assigned a corresponding crash likelihood. The GPS data from the MIT-AVT video clips was then mapped onto this heatmap so that each 20-meter segment is associated with an estimated crash likelihood. The resulting crash likelihood heatmap and the 500 video clips in the MIT-AVT dataset are shown in Fig. 4.
Figure 4: Crash likelihood heatmap (2018-2022) in Massachusetts, displayed in red, where darker colors indicate a higher crash likelihood. Five hundred video clips from the MIT-AVT dataset are marked in blue.
To keep the scale consistent with the Complexity Index, the crash likelihood is normalized and represented on a scale from 0 to 10. The 10,300 frames were categorized into three levels based on the calculated crash likelihood. As shown in Fig. 5, the crash likelihood distribution is skewed and unbalanced. There is a noticeable trend where crash likelihoods between 0 and 0.5 are highly frequent, decrease somewhat in the range of 0.5 to 2, and become increasingly rare beyond 2. Therefore, likelihoods between 0 and 0.5 were categorized as Low, those between 0.5 and 2 as Medium, and those between 2 and 10 as High. This classification resulted in approximately 3,600 images in the low category, 2,800 in the medium category, and 3,700 in the high category.
Figure 5: The distribution of crash likelihood.
@@ 3.4.1 Complexity-infused Encoder
The complexity-infused encoder uses a fully connected neural network structure with either 16 or 32 hidden neurons. The input to the network is threefold: (1) the 17 semantic features, (2) the combination of the 17 semantic features and 9 driving features, or (3) the combination of all features—17 semantic features, 9 driving features, and 19 contextual features. The input variables are normalized to a 0-1 range to ensure consistency across features and improve the stability of the model during training. The output of the network is the complexity index, which is treated as either a continuous or categorical variable for data obtained from the LLM, and as a continuous variable for data obtained from MTurk. The Root Mean Square Error (RMSE) is used as the evaluation metric for the complexity index when treated as a continuous variable, while accuracy is used as the metric when the complexity index is treated as a categorical variable.
@@
3.4.2 Crash Prediction Model
After generating the complexity-infused features from the encoder, they were used to predict the level of crash likelihood in combination with the corresponding input feature sets. For example, if the complexity-infused features were trained on only semantic features, the input for the crash prediction model would consist of both the complexity-infused and semantic features. Similarly, if the complexity-infused features were trained on all available features, the input to the crash prediction model would include the complexity-infused, semantic, driving, and contextual features.
To train the crash prediction model, various algorithms were tested, including Random Forest (RF), Gradient-Boosted Decision Trees (GBDT), K-Nearest Neighbors (KNN), and fully connected neural networks (NN). The NN model used in this step is a fully connected neural network consisting of seven linear layers. Similar to the encoder described in Section 3.4.1, the input variables are normalized to a 0-1 range. The dataset was split into 70% for training and 30% for testing, and model performance was evaluated using accuracy as the primary metric. To ensure consistency, the dataset was split in the same way as it was for the encoder.

### v2

# The Context of Crash Occurrence: A Complexity-Infused Approach Integrating Semantic, Contextual, and Kinematic Features

@@ Abstract
Understanding the context of crash occurrence in complex driving environments is essential for improving traffic safety and advancing automated driving. Previous studies have used statistical models and deep learning to predict crashes based on semantic, contextual, or vehicle kinematic features, but none have examined the combined influence of these factors. In this study, we term the integration of these features “roadway complexity”. This paper introduces a two-stage framework that integrates roadway complexity features for crash prediction. In the first stage, an encoder extracts hidden contextual information from these features, generating complexity-infused features. The second stage uses both original and complexity-infused features to predict crash likelihood, achieving an accuracy of 87.98% with original features alone and 90.15% with the added complexity-infused features. Ablation studies confirm that a combination of semantic, kinematic, and contextual features yields the best results, which emphasize their role in capturing roadway complexity. Additionally, complexity index annotations generated by the Large Language Model outperform those by Amazon Mechanical Turk, highlighting the potential of AI based tools for accurate, scalable crash prediction systems.
@@ 1 Introduction
Figure 1: The model structure. The model takes raw images and CAN-Bus signals as inputs to generate semantic, contextual, and kinematic features, which are then used to investigate their relationship with crash density estimates, serving as the model’s output. It consists of an encoder that learns hidden features from the semantic, kinematic, and contextual data, which are infused with the complexity index. The prediction model then utilizes all the available features, including the complexity-infused features, to predict the crash density and rates. Example data is shown above each feature source.
Understanding factors contributing to crashes is essential for improving traffic safety and advancing automated vehicle design. A 2015 NHTSA report [22] found that 94% of 5,470 crashes from 2005 to 2007 were due to human errors like inattention, distractions, excessive speed, and misjudgment.
Research has sought to predict crash frequency and rates based on environmental factors and driver behavior. Statistical models have linked crash occurrences to variables such as speed, traffic volume, weather, and road design [24, 21, 26, 7, 25], while recent studies have used deep learning to analyze real-time images and sensor data from vehicles and infrastructure to detect hazards [12, 9]. Though promising, these models rely on either imagery or behavioral data, without integrating both to predict crash occurrences or density.
@@
Driving behavior, such as speed adjustments in response to poor visibility or narrow lanes, is also influenced by scene complexity. Speed and acceleration patterns adjust based on obstacles and conditions [13]. Integrating behavior data with scene information deepens our understanding of driver interactions with their environment, improving crash risk modeling. Thus, we define roadway complexity as the combination of scene complexity and driving behavior, where driving behavior is represented by vehicle kinematic features.
Extracting hidden context from this combined data is essential. Previous studies have shown that fusing situational and memory-based features [29], as well as road networks and motion history data [18], enhances situation awareness and motion prediction, respectively. Building on this, we incorporate feature fusion to capture both explicit and implicit features of roadway scenes.
To address the challenges of (1) studying roadway complexity more holistically by incorporating both imagery and behavioral data, and (2) investigating the direct and indirect relationships between roadway complexity and crash density and rates, we propose a two-stage fusion prediction model framework. As illustrated in Fig. 1, the framework learns hidden features from the fusion of the semantic representation of the driving scene, vehicle kinematic features, and contextual features. The framework uses a combination of these features to form a comprehensive representation of roadway complexity and its association with crash density. Our model, which is trained on a naturalistic dataset and historical crash data, consists of an encoder that captures the hidden context of the roadway complexity and a prediction module that uses these features to investigate the relationship between the crash density. The contributions of this paper are:
@@
    We take a holistic approach to studying roadway complexity by considering both imagery information, i.e., specifically, semantic representation of the scene and contextual features, and vehicle kinematic data, and we empirically demonstrate the necessity of incorporating all these elements.
@@
    We argue that roadway complexity can be directly and indirectly linked to crash density and rates. We introduce a two-stage framework that extracts hidden contextual features of roadway scene complexity by fusing semantic, kinematic, and contextual data, and we investigate the predictive power of both direct and hidden features to identify features associated with crash density and rates.
@@
    We compare the complexity index annotations from Amazon Mechanical Turk and the Large Language Model (LLM) in terms of their capability to predict crash density and find that LLM-generated annotations consistently exhibited better predictive performance. This can enhance the development of real-time crash prediction systems and inspire the integration of automated annotation tools for improved accuracy and scalability.
@@ 2 Related Work
There has been extensive research on understanding roadway scenes using both Convolutional Neural Network (CNN)-based and transformer-based methods. Panoptic segmentation, which unifies semantic and instance segmentation, offers a holistic understanding by labeling all pixels and differentiating object instances, making it especially useful in structured environments like roadways [14]. The task has been successfully applied to driving scene datasets, such as CityScapes [3], which provides high-resolution images for semantic urban scene understanding. Recent advancements in transformer-based models have further improved performance in scene understanding. Among these, OneFormer [10] has achieved state-of-the-art panoptic quality scores [14] while maintaining computational efficiency, making it suitable for real-time scene understanding in automated driving. Given its strong performance, the OneFormer model was utilized in this study to extract scene-level visual features from the driving environments.
Understanding the driving context is essential for both human drivers and automated systems, as it involves not only the perception of roadway scene elements but also can help interpret the interactions between these elements. Recent advancements in LLMs have demonstrated their potential to go beyond pixel-level interpretation, incorporating higher-level reasoning into driving tasks. LLMs have been applied to extract contextual information from driving environments by mapping raw sensory data (e.g., images, LiDAR) to higher-level contextual descriptions [23]. This is particularly useful for driving research, where understanding the “intent” of other road users, predicting possible hazards, or inferring rules of the road from visual cues may significantly enhance safety and driver performance [4, 28]. The application of LLMs indicates a leap forward as they can go beyond object detection to understand the context of the driving scenarios.
@@ 3 Methods
In this section, we outline the dataset that was used in the study and the methodology for generating the features used in the crash density prediction model. We detail the architecture of the complexity-infused encoder and describe the structure of the final prediction model.
@@ 3.1 Dataset Overview
This study utilizes a subset dataset derived from the MIT-AVT naturalistic driving studies [6]. The dataset includes a variety of multimodal data sources, such as 20-second raw video clips from the forward-facing camera, 30 Hz CAN bus data, 30 Hz GPS data, and contextual metadata features. The contextual features are sourced from the study by Ding et al. [5].
For the purposes of this research, 500 video clips were selected from the dataset developed from Ding et al. [5], categorized as follows: 100 highway scenarios, 100 rural scenarios, 100 urban scenarios, 75 bridge scenarios, 75 overpass scenarios, and 75 crash hotspot scenarios. The methodology for generating crash hotspot data is detailed in Section 3.3.
From each selected video clip, frames were extracted based on a fixed-distance sampling method, with one frame captured for every 20 meters traveled. The extraction began with the first frame of each clip and continued until the distance to the subsequent frame was less than 20 meters. Overall, there were 10,407 frames extracted.
@@ 3.2 Feature Generation
For each image frame, three sets of features were generated: semantic features derived from a computer vision algorithm, vehicle kinematic features extracted from the CAN bus data, and contextual features generated from the LLM.
@@ 3.2.1 Semantic Features
The semantic features were generated using the OneFormer algorithm [10] via Hugging Face’s API [^1] https://huggingface.co/docs/transformers/en/model_doc/oneformer. The model outputs pixel-level semantic classifications for various objects such as cars, pedestrians, bicycles, roads, traffic signs, sidewalks, buildings, vegetation, and sky. To better understand the effect of complexity on driver behavior, a lead-car region was defined [27]. An example of the OneFormer algorithm’s output is provided in Fig. 2.
@@
For each image, the percentage of pixels corresponding to each class relative to the total frame size was calculated. Similarly, the percentage of pixels for each class within the lead-car region was computed. Additionally, the number of cars, pedestrians, buses, bicycles, and motorcycles was counted in both the full frame and the lead-car regions. This process resulted in 50 initial semantic features. After closer examination, features with minimal variability (more than $90\%$ being 00) were removed, reducing the final set to 17 semantic features. The final set of features is coded as:
@@
3.2.2 Vehicle Kinematic Features
The vehicle kinematic features were extracted from the CAN bus data. For each frame and the corresponding 20-meter segment, the following 9 features were computed: current speed, mean speed, standard deviation of speed, mean longitudinal acceleration, standard deviation of longitudinal acceleration, minimum longitudinal acceleration, maximum longitudinal acceleration, raw deviation from the speed limit, and normalized deviation from the speed limit. The raw and normalized deviation from the speed limit are defined as follows:
@@ 3.2.3 Contextual Features
The contextual features include road characteristics, which were generated using the GPT model [1]. The implementation details can be found in Appendix 7. The prompt used in this step is illustrated in Fig. 3. In the prompt, several questions were asked to gather contextual road characteristics, including information on weather conditions, road conditions, traffic conditions, visibility levels, time of day, road layout, road type, and lane width. Initially, each question was presented in an open-ended format to generate a pool of possible answers. The prompt was then refined based on these responses, providing predefined options to ensure consistency across the answers. During the prompting process, the GPT model was asked to generate a single output option for each question in the prompt, which was formatted in JSON. The prompt was run three times on each image to ensure data quality. For each question, the final answer was determined by selecting the response that was agreed upon by the majority of the three outputs.
@@ 3.3 Ground Truth Generation
There were two ground truths in this study. As illustrated in Figure 1, for the encoder, the complexity index was collected to quantify the overall complexity level in a given roadway scene. For the prediction, the crash density value was computed and serves as the ground truth for estimating crash rates.
@@ 3.3.1 Complexity Index
The complexity index was generated from two sources: AI and humans. For AI, the GPT model was used along with the contextual feature generation process, as shown in Fig. 3. In this approach, the model generated a complexity score on a scale from 0 to 10 to describe the complexity and demand level of the roadway scenes.
The human-generated complexity indices relied on Amazon Mechanical Turk (MTurk) for annotations. The task was designed to assess the complexity level of roadway scenes. Workers were shown image frames and asked to rate the complexity of each scene on a scale from 1 to 10. Only workers with a high approval rating, at least 500 completed tasks, and residing in the US were selected. A pilot study was conducted with 500 images, where 10 workers annotated the same image. The results showed a relatively high level of agreement among workers, with an average standard deviation of $1.74$ per image and a maximum standard deviation of less than $3$ for all images. Based on this, each scene for the study was annotated by 3 workers, and the final complexity score was determined by averaging their responses. Further details on the comparison between AI and human-generated annotations can be found in Appendix 8.
3.3.2 Crash Density
To generate the crash density, historical crash data was obtained from the Massachusetts Department of Transportation’s IMPACT app  [^3] https://apps.impact.dot.state.ma.us/cdp/home. GPS data from crashes over a 5-year period (2018 to 2022) was aggregated. A Kernel Density Estimation (KDE) method was then applied to this GPS data to create a continuous scale representing crash density. The Kernel Density Estimation is formulated as follows:
@@ 3.3.2 Crash Density
After applying the KDE method, each location was assigned a corresponding crash density value. The GPS data from the MIT-AVT video clips was then mapped onto this heatmap so that each 20-meter segment was associated with an estimated crash density. The resulting crash density heatmap and the 500 video clips selected from the MIT-AVT dataset are shown in Fig. 4.
Figure 4: Crash density heatmap (2018-2022) in Massachusetts, displayed in red, where darker colors indicate a higher crash density. Five hundred video clips from the MIT-AVT dataset are marked in blue.
To keep the scale consistent with the Complexity Index, the crash density was normalized and represented on a scale from 0 to 10. The 10,407 frames were categorized into three levels based on the calculated crash density value. As shown in Fig. 5, the crash density distribution was skewed and unbalanced. There was a noticeable trend where crash density ranges between 0 and 0.5 are highly frequent, decrease somewhat in the range of 0.5 to 2, and become increasingly rare beyond 2. Therefore, density values between 0 and 0.5 were categorized as Low, those between 0.5 and 2 as Medium, and those between 2 and 10 as High. This classification resulted in approximately 3,600 images in the low category, 2,800 in the medium category, and 3,700 in the high category.
Figure 5: The distribution of crash density value.
@@ 3.4.1 Complexity-infused Encoder
The complexity-infused encoder used a fully connected neural network structure with either 16 or 32 hidden neurons. The input to the network was threefold: (1) the 17 semantic features, (2) the combination of the 17 semantic features and 9 kinematic features, or (3) the combination of all features—17 semantic features, 9 kinematic features, and 19 contextual features. The input variables were normalized to a 0-1 range to ensure consistency across features and improve the stability of the model during training. The output of the network was the complexity index, which was treated as either a continuous or categorical variable for data obtained from the LLM, and as a continuous variable for data obtained from MTurk. The Root Mean Square Error (RMSE) was used as the evaluation metric for the complexity index when treated as a continuous variable, while accuracy was used as the metric when the complexity index was treated as a categorical variable.
@@
3.4.2 Crash Density Prediction Model
After generating the complexity-infused features from the encoder, they were used to predict the level of crash density value in combination with the corresponding input feature sets. For example, if the complexity-infused features were trained on only semantic features, the input for the crash prediction model would consist of both the complexity-infused and semantic features. Similarly, if the complexity-infused features were trained on all available features, the input to the crash prediction model would include the complexity-infused, semantic, kinematic, and contextual features.
To train the crash prediction model, various algorithms were tested, including Random Forest (RF), Gradient-Boosted Decision Trees (GBDT), K-Nearest Neighbors (KNN), and fully connected neural networks (NN). The NN model used in this step is a fully connected neural network consisting of seven linear layers. The detailed structure of the NN model can be found in Appendix 9. Similar to the encoder described in Section 3.4.1, the input variables were normalized to a 0-1 range. The dataset was split into 70% for training and 30% for testing, and model performance was evaluated using accuracy as the primary metric. To ensure consistency, the dataset was split in the same way as it was for the encoder.
For best-performing tree-based models, the SHAP values [15] were used to analyze the impact of the most influential features associated with crash density.